OpenStudy (anonymous):

Find the product of (x − 7)2

OpenStudy (aravindg):

apply distributive property a(b+c)=ab+ac

OpenStudy (anonymous):

ok thxs :)

OpenStudy (ajprincess):

is that 2 the exponent @Anna07?

OpenStudy (anonymous):

yeah

OpenStudy (ajprincess):

\((a-b)^2=a^2-2ab+b^2\)

OpenStudy (aravindg):

if its exponent then you need to use expansion above ..I though it was multiplication

OpenStudy (aravindg):

next time post like this to show exponent (x − 7)^2

OpenStudy (anonymous):

yeah I usually do.I don't know why I didn't this time